Balance Sheet
Brand Brew Inc.
Years 1 & 2 ($000's)
Year 1 . Year 2 .
Cash & Marketable
Securities 309,705 59,167
Accounts Receivable 108,732 705,426
Inventories 138,577 215,159
Other Current Assets 49,515 74,144
Total Current Assets 606,529 1,053,896
PP&E, Net 869,710 1,380,239
Intangibles 86,289 1,256,145
Other Assets 177,164 607,131
Total Assets 1,739,692 4,297,411
Accounts Payable 222,493 334,647
Other current liabilities 210,052 669,195
Short-term Debt 85,000 144,049
Total Current Liabilities 517,545 1,147,891
Long-term debt 20,000 1,383,392
Other long-term liabilities 250,835 784,277
Total liabilities 788,380 3,315,560
Capital Stock 8,922 28,334
Retained earnings 954,981 1,086,965
Adjustments -12,591 -133,448
Total shareholders' equity 951,312 981,851
Total Liabilities & Equity 1,739,692 4,297,411

Income Statement for Brand Brew Inc. Year 1 Year 2 Revenues 2,429,462 3,776,322 COGS 1,537,623 2,414,530 Depreciation 121,091 230,299 SG&A 619,143 833,208 EBIT 151,605 298,285 Interest Expense –14,403 49,732 Other income 32,005 8,047 Pre-Tax Income 198,013 256,600 Income Tax 75,049 94,947 Net Income 122,964 161,653

7. Refer to Brand Brew, Inc. financial statements,
Which of the following statements is correct?
(x) Both total asset turnover and ROA for the
firm decreased from Year 1 to Year 2.
(y) The amount of debt held by Brand Brew, Inc.
increased by more than 300 percent from
Year 1 to Year 2.
(z) The ROE for Brand Brew, Inc. rose from
12.9% in Year 1 to more than 17.4% in
Year 2.
A. (x), (y) and (z)
B. (x) and (y) only
C. (x) and (z) only
D. (y) and (z) only
E. (x) only


8. Refer to Brand Brew, Inc. financial statements.
What is the most important determinant of the
change in ROE?
A. ROA
B. Profit Margin
C. The change in leverage
D. Total Asset Turnover
E. Both A and B are important determinants
Answer #1

7. The correct answer is option B is only x and y are the correct statements.

working notes:

statement x

Total Asset Turnover = Net Sales / Average Total Assets

where Average total assets = total assets / 2

YEAR 1

Total Asset Turnover = 2429462 / 869846 = 2.792

Where, Average total Assets = 1739692 / 2 = 869846

Similarly in YEAR 2

Total Asset Turnover = 3776322 / 2148705 = 1.75

...

ROA = Net Income / Average total assets

YEAR 1

ROA = 122964 / 869846 = 0.141

YEAR 2

ROA = 161653 / 2148705 = 0.075

So clearly, X is true.

Statement y

Increase in total debt from year 1 to year 2 = 1527441 - 105000 = 1422441

percentage increase = 1422441 / 105000 = 13.547 * 100 = 1354.7%

Again, Y us true.

Statement z

ROE = Net income / Shareholders equity

YEAR 1, ROE = 122964 / 951312 = 0.12897 * 100 = 12.9 % (approx)

YEAR 2, ROE = 161653 / 981851 = 0.16464 * 100 = 16.4 %

Z is false because it says that roe in the year 2 is 17.4%.

.....

8. Option E ie both ROA And Profit margin.